OpenStudy (anonymous):

Solve 16^(d-4) = 3^(3-d) using logs

OpenStudy (mathstudent55):

First, take the log of both sides. Can you do that and show it?

OpenStudy (mathstudent55):

Here are the steps: 1. Take log of both sides. 2. Use the rule: \(\log a^b = b \log a\) on both sides to make the exponents become multipliers. 3. Distribute (d - 4) on the left side and (3 - d) on the right side. 4. Move all terms with d to the left side and all terms without d to the right side. 5. Factor out the d on the left side. 6. Divide both sides by the quantity multiplying d.

OpenStudy (mathstudent55):

I gave you the outline above. gtg. bye

OpenStudy (anonymous):

Thank You!

OpenStudy (mathstudent55):

You're welcome.